# Curbing PBHs with PTAs

### v1

@@ I Introduction
Primordial black holes (PBHs) Zel’dovich and Novikov (1967); Hawking (1971); Carr and Hawking (1974) are currently in the spotlight as they may solve several open questions in astrophysics and cosmology. Different mechanisms and models have been proposed as viable for PBH production and, depending on the formation scenario considered and on the models for PBH formations, a variety of mass functions have been predicted (see Ref. Bagui et al. (2023) for a recent review). Throughout this work, we will assume the standard formation scenario which suggests that PBHs form out of the gravitational collapse after the horizon re-entry of large over-densities in the primordial density contrast field Ivanov et al. (1994); Ivanov (1998); Blinnikov et al. (2016). In such a scenario, a non-negligible PBH abundance requires an enhanced primordial curvature power spectrum at scales smaller than the scale of the Cosmic Microwave Background (CMB), at which the spectral amplitude is around $10^{-9}$ Akrami et al. (2020a).
@@
In addition to producing PBHs, the same primordial curvature perturbations generate tensor perturbations, also known as scalar-induced Gravitational Waves (SIGW). As first shown by Refs. Matarrese et al. (1994); Acquaviva et al. (2003); Mollerach et al. (2004); Ananda et al. (2007); Baumann et al. (2007), when going to second order, scalar and tensor perturbations are coupled and the former provide a source for Gravitational Waves (GW) (see also Refs. Espinosa et al. (2018); Domènech (2021)). The statistical properties of the curvature fluctuations can leave an imprint on this signal. Indeed, the presence of primordial non-Gaussianity (NG) Gangui et al. (1994); Bartolo et al. (2002); Maldacena (2003); Bartolo et al. (2004a, b); Celoria and Matarrese (2020) can modify the spectrum inducing specific features depending on the amplitude of the sourcing scalar power spectrum Garcia-Saenz et al. (2023); Unal (2019); Adshead et al. (2021); Abe et al. (2023); Perna et al. (2024). Moreover, the specific features of the spectra and the NG modifications, as well as the amplitude of the induced SIGW background, depend not only on the shape of the curvature power spectrum considered, but also on the specific scales involved (see, e.g., Cai et al. (2019a, b)). Hence detectable SIGWs offer a potential way to probe early universe cosmology.
@@
The recent Pulsar Timing Array (PTA) data release by the NANOGrav Agazie et al. (2023a, b), EPTA (in combination with InPTA) Antoniadis et al. (2023a, b, c), PPTA Reardon et al. (2023a); Zic et al. (2023); Reardon et al. (2023b) and CPTA Xu et al. (2023) collaborations shows evidence of a Hellings-Downs pattern in the angular correlations which is characteristic of GWs. The most stringent constraints and largest statistical evidence come from the NANOGrav 15-year data (NANOGrav15). Even though it is currently not possible to pinpoint the sources of this signal, this discovery is of extreme importance since it opens a further window into the early universe and has an impact on the GW search at the nHz scale (see, e.g., Ellis et al. (2024); Figueroa et al. (2024)).
@@
Since PTAs are sensitive to frequencies of the order of the nHz, they are associated with the production of PBHs in the stellar mass range Chen et al. (2020); De Luca et al. (2021); Vaskonen and Veermäe (2021); Kohri and Terada (2021); Domènech and Pi (2022); Dandoy et al. (2023); Franciolini et al. (2023a). Crucially, requiring that the produced SIGW background does not exceed the one registered by the NANOGrav15, strongly limits the maximum amplitude of the curvature power spectrum, and thereby the abundance of PBHs in the corresponding mass range. Additionally, if the power spectrum is enhanced at scales larger than accessible by PTAs, PBHs can serve as seeds of supermassive black holes (SMBHs) Duechting (2004); Bernal et al. (2018); Serpico et al. (2020). Nevertheless, in this case, the scales related to the formation process are strongly constrained by CMB spectral distortion analysis of the FIRAS collaboration Fixsen et al. (1996); Chluba et al. (2012a, b); Chluba and Grin (2013); Bianchini and Fabbian (2022) and, without NGs, the allowed PBH abundance is too small to provide a primordial origin for SMBHs seeds.
@@
In this work constraints on PBHs arising from NANOGrav15 are considered in detail with an emphasis on local primordial NGs and assuming different shapes for the primordial power spectrum. After having reviewed the origin of NGs in the standard formation scenario for PBHs in Sec. II, in Sec. III we discuss different formalisms for the computation of the PBH abundance and the role of NGs. In Sec. IV we summarise the computation of the SIGW background. Then, in Sec. V, we derive PTA constraints on the power spectrum and the abundance of PBHs, and comment on how large positive NGs affect the generation of PBHs as seeds for SMBHs. We conclude in Sec. VI.
II Non-Gaussian primordial curvature perturbations
@@
$\mathcal{P}_{\zeta}(k)=\frac{{A}}{\sqrt{2\pi}\Delta}\exp\left(-\frac{1}{2\Delta^{2}}\ln^{2}\left(\frac{k}{k_{*}}\right)\right)\,,$ (1)
@@
$\mathcal{P}_{\zeta}(k)={A}\,\frac{\left(\alpha+\beta\right)^{\lambda}}{\left[\beta\left(k/k_{*}\right)^{-\alpha/\lambda}+\alpha\left(k/k_{*}\right)^{\beta/\lambda}\right]^{\lambda}},$ (2)
where the parameters $\alpha,\beta>0$ describe the growth and decay of the spectrum around the peak and $\lambda$ the width of the peak. Typically, $\alpha\approx 4$ Byrnes et al. (2019); Karam et al. (2023a) and $\beta\lesssim 4$. The cases considered in this work are reported in Tab. 1.
@@
LN LN                                    $(\Delta=0.5)$ $[0,10]$
BPL1 BPL                  $(\alpha=4,\beta=3,\lambda=1)$ $[0,10]$
BPL2 BPL                  $(\alpha=4,\beta=0.5,\lambda=1)$ $[0,10]$
@@
The standard formation scenario assumes that PBHs arise from the gravitational collapse of significant over-densities in the primordial density contrast field Ivanov et al. (1994); Ivanov (1998). Given the random nature of this field, the calculation of PBH abundance is inherently statistical and requires precise knowledge of the probability density function (PDF) of the density fluctuations.
@@
Limiting the analysis to the assumption that the PDF of density fluctuations $\delta$ follows the Gaussian statistics is theoretically flawed for two reasons. First, density fluctuations in the primordial radiation field originate from curvature perturbations $\zeta$, previously stretched on super-horizon scales during inflation, after their horizon re-entry. In the long-wavelength approximation, the equation which relates curvature perturbations to density fluctuations is intrinsically non-linear Harada et al. (2015)
$\delta(\vec{x},t)\!=\!-\frac{2}{3}\frac{\Phi}{(aH)^{2}}e^{-2\zeta(\vec{x})}\bigg{[}\nabla^{2}\zeta(\vec{x})\!+\!\frac{1}{2}\partial_{i}\zeta(\vec{x})\partial_{i}\zeta(\vec{x})\bigg{]},\!\!$ (3)
@@
where $a$ denotes the scale factor, $H$ the Hubble rate, $\Phi$ is related to the equation of state parameter ${w_{\rm u}}$ of the universe. For ${w_{\rm u}}$ constant, $\Phi=3(1+{w_{\rm u}})/(5+3{w_{\rm u}})$ Polnarev and Musco (2007). We dropped the explicit $\vec{x}$ and $t$ dependence for brevity. For this reason, even under the assumption that curvature perturbations follow exact Gaussian statistics, density fluctuations inherit an unavoidable amount of NG from non-linear (NL) corrections De Luca et al. (2019); Young et al. (2019); Germani and Sheth (2020). Second, there is no guarantee that $\zeta$ is a Gaussian field – we refer to such cases as primordial NGs. Local primordial NGs can be described by a relation
$\zeta(\vec{x})=F(\zeta_{\rm G}(\vec{x}))$ (4)
@@
between $\zeta$ and its Gaussian counterpart $\zeta_{G}$, which depends on the mechanism that generates the enhanced power spectrum at small scales (see the related discussion in Ref. Ferrante et al. (2023a)). Such NGs are thus generically independent of large-scale NGs constrained by CMB data (e.g. Akrami et al. (2020b)). Copious PBH production via critical collapse of large overdensities is achievable in a wide range of scenarios including single-field inflation with an inflation point in the inflaton’s potential (usually called USR models) Garcia-Bellido and Ruiz Morales (2017); Pi et al. (2018); Kannike et al. (2017); Ballesteros et al. (2020); Inomata et al. (2017); Iacconi et al. (2022); Kawai and Kim (2021); Bhaumik and Jain (2020); Cheong et al. (2021); Inomata et al. (2018); Dalianis et al. (2019); Motohashi et al. (2020); Hertzberg and Yamada (2018); Ballesteros and Taoso (2018); Karam et al. (2023a); Rasanen and Tomberg (2019); Balaji et al. (2022); Frolovsky and Ketov (2023); Dimopoulos (2017); Germani and Prokopec (2017); Choudhury and Mazumdar (2014); Ragavendra and Sriramkumar (2023); Cheng et al. (2022); Franciolini et al. (2023b); Karam et al. (2023b); Mishra et al. (2023); Cole et al. (2023); Karam et al. (2023b); Frosina and Urbano (2023); Franciolini and Urbano (2022); Choudhury et al. (2024); Wang et al. (2024); Stamou (2021, 2024); Heydari and Karami (2022a, b, 2024) and models with spectator field, i.e., the curvaton Enqvist and Sloth (2002); Lyth and Wands (2002); Sloth (2003); Lyth et al. (2003); Dimopoulos et al. (2003); Kohri et al. (2013); Kawasaki et al. (2013a, b); Carr et al. (2017); Ando et al. (2018a, b); Chen and Cai (2019); Liu and Prokopec (2021); Pi and Sasaki (2023); Cai et al. (2021); Liu (2023); Chen et al. (2023); Torrado et al. (2018); Chen et al. (2023); Wilkins and Cable (2024); Inomata et al. (2024).
@@
$\zeta=\zeta_{\rm G}+\frac{3}{5}f_{\rm NL}(\zeta_{\rm G}^{2}-\langle\zeta_{\rm G}^{2}\rangle)+\frac{9}{25}g_{\rm NL}\zeta_{\rm G}^{3}+\dots\,,$ (5)
@@ III PBH formation
$M_{\rm PBH}=\mathcal{K}M_{k}(\mathcal{C}-\mathcal{C}_{\rm th})^{\gamma}\,,$ (6)
@@
$M_{k}\approx 17M_{\odot}\left(\frac{k}{10^{6}{\rm Mpc}^{-1}}\right)^{-2}\,\left(\frac{g_{*}}{10.75}\right)^{-1/6}\,$ (7)
is the mass contained within a Hubble horizon corresponding to a comoving scale $k$. The threshold $\mathcal{C}_{\rm th}$ is estimated following Ref. Musco et al. (2021) and depends on the shape of the power spectrum. We also account for the time dependence of $\mathcal{C}_{\rm th}$, the critical exponent $\gamma$ and the prefactor $\mathcal{K}$ during the QCD phase transition Musco et al. (2024).
@@
$\!\!\!\beta_{k}(M_{\rm PBH})\!=\!\int_{\mathcal{C}_{\rm th}}^{\infty}\!{\rm d}\mathcal{C}\frac{M_{\rm PBH}(\mathcal{C})}{M_{k}}P_{k}(\mathcal{C})\delta\left[\ln\frac{M_{\rm PBH}}{M_{\rm PBH}(\mathcal{C})}\right]\!,$ (8)
@@
$\!\frac{{\rm d}f_{\rm PBH}}{{\rm d}\ln M_{\rm PBH}}$ $\!=\!\frac{1}{\Omega_{\rm DM}}\int\frac{{\rm d}M_{k}}{M_{k}}\,\beta_{k}(M_{\rm PBH})\left(\frac{M_{\rm eq}}{M_{k}}\right)^{1/2}\!\!,$ (9)
where $M_{\rm eq}\approx 2.8\times 10^{17}\,\,M_{\odot}$ is the horizon mass at the time of matter-radiation equality and $\Omega_{\rm DM}=0.12h^{-2}$ is the cold dark matter density Akrami et al. (2020a). To characterise the PBH population, we will consider the PBH abundance and the mean PBH mass,
$f_{\rm PBH}$ $=\int\frac{{\rm d}M_{\rm PBH}}{M_{\rm PBH}}\frac{{\rm d}f_{\rm PBH}}{{\rm d}\ln M_{\rm PBH}}\,,$ (10)
@@
where the mean PBH mass is computed with respect to the PBH number density (see e.g. Andrés-Carcasona et al. (2024)). This offers a clear basis for comparing constraints on PBHs with extended mass functions.
@@
dictates that the PBH formation probability can be estimated from the statistics of the compaction function $\mathcal{C}$ Ferrante et al. (2023a); Gow et al. (2023), generically defined as twice the local mass excess over the areal radius [^1] Using the threshold statistics approach, which relies on average compaction profiles Ianniccari et al. (2024), corrections to the horizon crossing and from the non-linear radiation transfer function Franciolini et al. (2023c); De Luca et al. (2023) can affect the PBH abundance. Since the amount of these corrections is still not well understood we leave their inclusions for a future work.. The latter can be estimated in scenarios with non-Gaussian curvature perturbations $\zeta$ in case they can be characterized via an auxiliary Gaussian field $\zeta_{G}$ as outlined in Sec. II. The compaction function $\mathcal{C}=\mathcal{C}_{1}-\mathcal{C}_{1}^{2}/(4\Phi)$ can then be constructed from $\mathcal{C}_{1}=\mathcal{C}_{\rm G}\,{\rm d}F/{\rm d}\zeta_{\rm G}$ which can be expressed in terms of the Gaussian field $\mathcal{C}_{\rm G}=-2\Phi\,r\,\zeta_{\rm G}^{\prime}$.
@@
$P_{k}(\mathcal{C})=\int_{\mathcal{D}}\delta(\mathcal{C}-\mathcal{C}(\mathcal{C}_{\rm G},\zeta_{\rm G}))P_{{\rm G},k}(\mathcal{C}_{\rm G},\zeta_{\rm G}){\rm d}\mathcal{C}_{\rm G}{\rm d}\zeta_{\rm G}\,,$ (11)
@@
$P_{\mathrm{G}}\left(\mathcal{C}_{\mathrm{G}},\zeta_{\mathrm{G}}\right)=\frac{e^{\left[-\frac{1}{2\left(1-\gamma_{cr}^{2}\right)}\left(\frac{\mathcal{C}_{\mathrm{G}}}{\sigma_{c}}-\frac{\gamma_{cr}\zeta_{\mathrm{G}}}{\sigma_{r}}\right)^{2}\!-\!\frac{\zeta_{\mathrm{G}}^{2}}{2\sigma_{r}^{2}}\right]}}{2\pi\sigma_{c}\sigma_{\tau}\sqrt{1-\gamma_{cr}^{2}}}.$ (12)
@@
$\sigma_{c}^{2}=\frac{4\Phi^{2}}{9}\int_{0}^{\infty}\frac{{\rm d}k}{k}\left(kr_{m}\right)^{4}W^{2}\left(k,r_{m}\right)P^{T}_{\zeta}\,,$ (13a)
$\sigma_{cr}^{2}=\frac{2\Phi}{3}\!\int_{0}^{\infty}\!\!\frac{{\rm d}k}{k}\!\left(kr_{m}\right)^{2}\!W\!\!\left(k,r_{m}\right)\!W_{s}\!\left(k,r_{m}\right)\!P^{T}_{\zeta}\!\,,$ (13b)
$\sigma_{r}^{2}=\int_{0}^{\infty}\frac{{\rm d}k}{k}W_{s}^{2}\left(k,r_{m}\right)P^{T}_{\zeta}\,,$ (13c)
with $P^{T}_{\zeta}=T^{2}\left(k,r_{m}\right)P_{\zeta}(k)$, and $\gamma_{cr}\equiv\sigma_{cr}^{2}/\sigma_{c}\sigma_{\tau}$. We have defined $W\left(k,r_{m}\right),$ $W_{s}\left(k,r_{m}\right)$ and $T\left(k,r_{m}\right)$ as the top-hat window function, the spherical-shell window function, and the radiation transfer function Young (2022).
The PBH mass function (9) can be expressed as
@@
×$\times\frac{{\cal K}}{\gamma\sqrt{\Lambda}}\int{\rm d}\zeta_{{\rm G},k}P_{\rm G}({\cal C}_{\rm G}(M_{\rm PBH},\zeta_{\rm G}),\zeta_{\rm G})\left(\frac{dF}{d\zeta_{\rm G}}\right)^{-1}\!\!\!\!,$ (14)
@@
It is well known that the Press-Schechter approach does not agree with the theory of peaks (see, e.g., Refs. Green et al. (2004); Young et al. (2014); De Luca et al. (2019)). A technical drawback of peaks theory is that it is not clear how to include NGs in the computation of the abundance in this approach using a generic functional form for the curvature perturbation field as in Eq. (4). Hence, for comparison, we limit our analysis for the peak theory approach to the case of negligible primordial NGs [^2] The impact of local-type NG on PBH abundance in peak theory is expected to be reduced, especially for higher-order terms Young (2022) and for peaked power spectra. Indeed, the compaction is volume-averaged over the scale of the perturbation and depends on the curvature perturbation at its edge, rather than its centre..
In this approach, the starting point is the number density of peaks with a height in the interval $(\nu,\nu+{\rm d}\nu)$. When $\nu\gg 1$, it is given by Bardeen et al. (1986)
$\frac{{\rm d}\mathcal{N}}{{\rm d}\nu}=\frac{1}{4\pi^{2}r_{m}^{3}}\left(\frac{\sigma_{cc}}{\sigma_{c}}\right)^{3}\nu^{3}\exp\left(-\frac{\nu^{2}}{2}\right)\,,$ (15)
where we introduced the rescaled peak height $\nu\equiv\mathcal{C}_{1}/\sigma_{c}$ and the first rescaled moment of the distribution Balaji et al. (2023)
$\sigma_{cc}^{2}=\frac{4\Phi^{2}}{9}\int_{0}^{\infty}\frac{{\rm d}k}{k}\left(kr_{m}\right)^{6}W^{2}\left(k,r_{m}\right)P^{T}_{\zeta}\,.$ (16)

### v2

@@ 1 Introduction
Primordial black holes (PBHs) [1, 2, 3] are currently in the spotlight as they may solve several open questions in astrophysics and cosmology. Different mechanisms and models have been proposed as viable for PBH production and, depending on the formation scenario considered and on the models for PBH formations, a variety of mass functions have been predicted (see Ref. [4] for a recent review). Throughout this work, we will assume the standard formation scenario which suggests that PBHs form out of the gravitational collapse after the horizon re-entry of large over-densities in the primordial density contrast field [5, 6, 7]. In such a scenario, a non-negligible PBH abundance requires an enhanced primordial curvature power spectrum at scales smaller than the scale of the Cosmic Microwave Background (CMB), at which the spectral amplitude is around $10^{-9}$ [8].
@@
In addition to producing PBHs, the same primordial curvature perturbations generate tensor perturbations, also known as scalar-induced Gravitational Waves (SIGW). As first shown by Refs. [9, 10, 11, 12, 13], when going to second order, scalar and tensor perturbations are coupled and the former provide a source for Gravitational Waves (GW) (see also Refs. [14, 15]). The statistical properties of the curvature fluctuations can leave an imprint on this signal. Indeed, the presence of primordial non-Gaussianity (NG) [16, 17, 18, 19, 20, 21] can modify the spectrum inducing specific features depending on the amplitude of the sourcing scalar power spectrum [22, 23, 24, 25, 26]. Moreover, the specific features of the spectra and the NG modifications, as well as the amplitude of the induced SIGW background, depend not only on the shape of the curvature power spectrum considered, but also on the specific scales involved (see, e.g., [27, 28]). Hence detectable SIGWs offer a potential way to probe early universe cosmology.
@@
The recent Pulsar Timing Array (PTA) data release by the NANOGrav [29, 30], EPTA (in combination with InPTA) [31, 32, 33], PPTA [34, 35, 36] and CPTA [37] collaborations shows evidence of a Hellings-Downs pattern in the angular correlations which is characteristic of GWs. The most stringent constraints and largest statistical evidence come from the NANOGrav 15-year data (NANOGrav15). Even though it is currently not possible to pinpoint the sources of this signal, this discovery is of extreme importance since it opens a further window into the early universe and has an impact on the GW search at the nHz scale (see, e.g., [38, 39]).
@@
Since PTAs are sensitive to frequencies of the order of the nHz, they are associated with the production of PBHs in the stellar mass range [40, 41, 42, 43, 44, 45, 46, 47]. Crucially, requiring that the produced SIGW background does not exceed the one registered by the NANOGrav15, strongly limits the maximum amplitude of the curvature power spectrum, and thereby the abundance of PBHs in the corresponding mass range. Additionally, if the power spectrum is enhanced at scales larger than accessible by PTAs, PBHs can serve as seeds of supermassive black holes (SMBHs) [48, 49, 50]. Nevertheless, in this case, the scales related to the formation process are strongly constrained by CMB spectral distortion analysis of the FIRAS collaboration [51, 52, 53, 54, 55] and, without NGs, the allowed PBH abundance is too small to provide a primordial origin for SMBHs seeds.
@@
In this work constraints on PBHs arising from NANOGrav15 are considered in detail with an emphasis on local primordial NGs and assuming different shapes for the primordial power spectrum. After reviewing the origin of NGs in the standard formation scenario for PBHs in Sec. 2, in Sec. 3 we discuss different formalisms for estimating the PBH abundance and the role of NGs. In Sec. 4 we summarise the computation of the SIGW background. Then, in Sec. 5, we derive PTA constraints on the power spectrum and the abundance of PBHs, and comment on how large positive NGs affect the generation of PBHs as seeds for SMBHs. We conclude in Sec. 6.
2 Non-Gaussian primordial curvature perturbations
@@
$\mathcal{P}_{\zeta}(k)=\frac{{A}}{\sqrt{2\pi}\Delta}\exp\left(-\frac{1}{2\Delta^{2}}\ln^{2}\left(\frac{k}{k_{*}}\right)\right)\,,$ (2.1)
@@
$\mathcal{P}_{\zeta}(k)={A}\,\frac{\left(\alpha+\beta\right)^{\lambda}}{\left[\beta\left(k/k_{*}\right)^{-\alpha/\lambda}+\alpha\left(k/k_{*}\right)^{\beta/\lambda}\right]^{\lambda}},$ (2.2)
where the parameters $\alpha,\beta>0$ describe the growth and decay of the spectrum around the peak and $\lambda$ the width of the peak. Typically, $\alpha\approx 4$ [56, 57] and $\beta\lesssim 4$. The cases considered in this work are reported in Tab. 1.
@@
LN LN $(\Delta=0.5)$ $[-2,10]$
BPL1 BPL $(\alpha=4,\beta=3,\lambda=1)$ $[-2,10]$
BPL2 BPL $(\alpha=4,\beta=0.5,\lambda=1)$ $[-2,10]$
@@
The standard formation scenario assumes that PBHs arise from the gravitational collapse of significant over-densities in the primordial density contrast field [5, 6]. Given the random nature of this field, the calculation of PBH abundance is inherently statistical and requires precise knowledge of the probability density function (PDF) of the density fluctuations.
@@
Limiting the analysis to the assumption that the PDF of density fluctuations $\delta$ follows the Gaussian statistics is theoretically flawed for two reasons. First, density fluctuations in the primordial radiation field originate from curvature perturbations $\zeta$, previously stretched on super-horizon scales during inflation, after their horizon re-entry. In the long-wavelength approximation, the equation which relates curvature perturbations to density fluctuations is intrinsically non-linear [58]
$\delta(\vec{x},t)\!=\!-\frac{2}{3}\frac{\Phi}{(aH)^{2}}e^{-2\zeta(\vec{x})}\bigg{[}\nabla^{2}\zeta(\vec{x})\!+\!\frac{1}{2}\partial_{i}\zeta(\vec{x})\partial_{i}\zeta(\vec{x})\bigg{]},\!\!$ (2.3)
@@
where $a$ denotes the scale factor, $H$ the Hubble rate, $\Phi$ is related to the equation of state parameter ${w_{\rm u}}$ of the universe. For ${w_{\rm u}}$ constant, $\Phi=3(1+{w_{\rm u}})/(5+3{w_{\rm u}})$ [59]. We dropped the explicit $\vec{x}$ and $t$ dependence for brevity. For this reason, even under the assumption that curvature perturbations follow exact Gaussian statistics, density fluctuations inherit an unavoidable amount of NG from non-linear (NL) corrections [60, 61, 62]. Second, there is no guarantee that $\zeta$ is a Gaussian field – we refer to such cases as primordial NGs. Local primordial NGs can be described by a relation
$\zeta(\vec{x})=F(\zeta_{\rm G}(\vec{x}))$ (2.4)
@@
between $\zeta$ and its Gaussian counterpart $\zeta_{G}$, which depends on the mechanism that generates the enhanced power spectrum at small scales (see the related discussion in Ref. [63]). Such NGs are thus generically independent of large-scale NGs constrained by CMB data (e.g. [64]). Copious PBH production via critical collapse of large overdensities is achievable in a wide range of scenarios including single-field inflation with an inflation point in the inflaton’s potential (usually called USR models) [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 57, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 88, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100] and models with spectator field, i.e., the curvaton [101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 117, 119, 120].
@@
$\zeta=\zeta_{\rm G}+\frac{3}{5}f_{\rm NL}(\zeta_{\rm G}^{2}-\langle\zeta_{\rm G}^{2}\rangle)+\frac{9}{25}g_{\rm NL}\zeta_{\rm G}^{3}+\dots\,,$ (2.5)
@@ 3 PBH formation
$M_{\rm PBH}=\mathcal{K}M_{k}(\mathcal{C}-\mathcal{C}_{\rm th})^{\gamma}\,,$ (3.1)
@@
$M_{k}\approx 17M_{\odot}\left(\frac{k}{10^{6}{\rm Mpc}^{-1}}\right)^{-2}\,\left(\frac{g_{*}}{10.75}\right)^{-1/6}\,$ (3.2)
is the mass contained within a Hubble horizon corresponding to a comoving scale $k$. The threshold $\mathcal{C}_{\rm th}$ is estimated following Ref. [121] and depends on the shape of the power spectrum $q$. In Fig. 1 we show the variation of the threshold as a function of the horizon mass due to the QCD phase transition and its dependence on the spectral shape parameter $q$.
Figure 1: Left panel: Threshold $C_{\rm th}$ as a function of the shape parameter $q$. The coloured dots indicate the three benchmark scenarios in Table 1. Right panel: Relative change of threshold $C_{\rm th}$ as a function of the horizon mass $M_{\rm H}$.
We also account for the time dependence of $\mathcal{C}_{\rm th}$, the critical exponent $\gamma$ and the prefactor $\mathcal{K}$ during the QCD phase transition [122]. We fix $\mathcal{K}=4.4$ and $\gamma=0.38$.
@@
$\beta_{k}(M_{\rm PBH})=\int_{\mathcal{C}_{\rm th}}\!{\rm d}\mathcal{C}\,P_{k}(\mathcal{C})\frac{M_{\rm PBH}}{M_{k}}\delta\left[\ln\frac{M_{\rm PBH}}{M_{\rm PBH}(\mathcal{C})}\right]\!,$ (3.3)
@@
$\!\frac{{\rm d}f_{\rm PBH}}{{\rm d}\ln M_{\rm PBH}}$ $\!=\!\frac{1}{\Omega_{\rm DM}}\int\frac{{\rm d}M_{k}}{M_{k}}\,\beta_{k}(M_{\rm PBH})\left(\frac{M_{\rm eq}}{M_{k}}\right)^{1/2}\!\!,$ (3.4)
where $M_{\rm eq}\approx 2.8\times 10^{17}\,\,M_{\odot}$ is the horizon mass at the time of matter-radiation equality and $\Omega_{\rm DM}=0.12h^{-2}$ is the cold dark matter density [8]. To characterise the PBH population, we will consider the PBH abundance and the mean PBH mass,
$f_{\rm PBH}$ $=\int\frac{{\rm d}M_{\rm PBH}}{M_{\rm PBH}}\frac{{\rm d}f_{\rm PBH}}{{\rm d}\ln M_{\rm PBH}}\,,$ (3.5)
@@
where the mean PBH mass is computed with respect to the PBH number density (see e.g. [123]). This offers a clear basis for comparing constraints on PBHs with extended mass functions.
@@
dictates that the PBH formation probability can be estimated from the statistics of the compaction function $\mathcal{C}$ [63, 124], generically defined as twice the local mass excess over the areal radius [^1] Using the threshold statistics approach, which relies on average compaction profiles [125], corrections to the horizon crossing and from the non-linear radiation transfer function [126, 127] can affect the PBH abundance. Since the amount of these corrections is still not well understood we leave their inclusions for future work.. The latter can be estimated in scenarios with NG curvature perturbations $\zeta$ in case they can be characterized via an auxiliary Gaussian field $\zeta_{G}$ as outlined in Sec. 2. The compaction function $\mathcal{C}=\mathcal{C}_{1}-\mathcal{C}_{1}^{2}/(4\Phi)$ can then be constructed from $\mathcal{C}_{1}=\mathcal{C}_{\rm G}\,{\rm d}F/{\rm d}\zeta_{\rm G}$ which can be expressed in terms of the Gaussian field $\mathcal{C}_{\rm G}=-2\Phi\,r\,\zeta_{\rm G}^{\prime}$.
@@
$P_{k}(\mathcal{C})=\int_{\mathcal{D}}\delta(\mathcal{C}-\mathcal{C}(\mathcal{C}_{\rm G},\zeta_{\rm G}))P_{{\rm G},k}(\mathcal{C}_{\rm G},\zeta_{\rm G}){\rm d}\mathcal{C}_{\rm G}{\rm d}\zeta_{\rm G}\,,$ (3.6)
@@
$P_{\mathrm{G}}\left(\mathcal{C}_{\mathrm{G}},\zeta_{\mathrm{G}}\right)=\frac{e^{\left[-\frac{1}{2\left(1-\gamma_{cr}^{2}\right)}\left(\frac{\mathcal{C}_{\mathrm{G}}}{\sigma_{c}}-\frac{\gamma_{cr}\zeta_{\mathrm{G}}}{\sigma_{r}}\right)^{2}\!-\!\frac{\zeta_{\mathrm{G}}^{2}}{2\sigma_{r}^{2}}\right]}}{2\pi\sigma_{c}\sigma_{\tau}\sqrt{1-\gamma_{cr}^{2}}}.$ (3.7)
@@
$\sigma_{c}^{2}=\frac{4\Phi^{2}}{9}\int_{0}^{\infty}\frac{{\rm d}k}{k}\left(kr_{m}\right)^{4}W^{2}\left(k,r_{m}\right)P^{T}_{\zeta}\,,$ (3.8a)
$\sigma_{cr}^{2}=\frac{2\Phi}{3}\!\int_{0}^{\infty}\!\!\frac{{\rm d}k}{k}\!\left(kr_{m}\right)^{2}\!W\!\!\left(k,r_{m}\right)\!W_{s}\!\left(k,r_{m}\right)\!P^{T}_{\zeta}\!\,,$ (3.8b)
$\sigma_{r}^{2}=\int_{0}^{\infty}\frac{{\rm d}k}{k}W_{s}^{2}\left(k,r_{m}\right)P^{T}_{\zeta}\,,$ (3.8c)
with $P^{T}_{\zeta}=T^{2}\left(k,r_{m}\right)P_{\zeta}(k)$, and $\gamma_{cr}\equiv\sigma_{cr}^{2}/\sigma_{c}\sigma_{\tau}$. We have defined $W\left(k,r_{m}\right),$ $W_{s}\left(k,r_{m}\right)$ and $T\left(k,r_{m}\right)$ as the top-hat window function, the spherical-shell window function, and the radiation transfer function [128] [^2] The choice of the window function is dictated by the approach used to evaluate the threshold for the critical collapse. Using different window functions is possible, but it would be necessary to reevaluate the threshold in order to maintain the accuracy of the evaluation [129]..
The PBH mass function (3.4) can be expressed as
@@
×$\times\frac{{\cal K}}{\gamma\sqrt{\Lambda}}\int{\rm d}\zeta_{{\rm G},k}P_{\rm G}({\cal C}_{\rm G}(M_{\rm PBH},\zeta_{\rm G}),\zeta_{\rm G})\left(\frac{dF}{d\zeta_{\rm G}}\right)^{-1}\!\!\!\!,$ (3.9)
@@
It is well known that the Press-Schechter approach does not agree with the theory of peaks (see, e.g., Refs. [130, 131, 60]). A technical drawback of peaks theory is that it is not clear how to include NGs in the computation of the abundance in this approach using a generic functional form for the curvature perturbation field as in Eq. (2.4). Hence, for comparison, we limit our analysis for the peak theory approach to the case of negligible primordial NGs [^3] The impact of local-type NG on PBH abundance in peak theory is expected to be reduced, especially for higher-order terms [128] and for peaked power spectra. Indeed, the compaction is volume-averaged over the scale of the perturbation and depends on the curvature perturbation at its edge, rather than its centre..
In this approach, the starting point is the number density of peaks with a height in the interval $(\nu,\nu+{\rm d}\nu)$. When $\nu\gg 1$, it is given by [132]
$\frac{{\rm d}\mathcal{N}}{{\rm d}\nu}=\frac{1}{4\pi^{2}r_{m}^{3}}\left(\frac{\sigma_{cc}}{\sigma_{c}}\right)^{3}\nu^{3}\exp\left(-\frac{\nu^{2}}{2}\right)\,,$ (3.10)
where we introduced the rescaled peak height $\nu\equiv\mathcal{C}_{1}/\sigma_{c}$ and the first rescaled moment of the distribution [133]
$\sigma_{cc}^{2}=\frac{4\Phi^{2}}{9}\int_{0}^{\infty}\frac{{\rm d}k}{k}\left(kr_{m}\right)^{6}W^{2}\left(k,r_{m}\right)P^{T}_{\zeta}\,.$ (3.11)